OpenStudy (ashontae19):

The height of a triangular window is (x + 7) cm and the base measures (3x) cm. If the area of the window is 66 cm2, what is the height of the window?

OpenStudy (ashontae19):

A. 12 cm B. 11 cm C. 51 cm D. 4 cm

OpenStudy (ashontae19):

can you help me?

OpenStudy (isaiah.feynman):

What you can do is to use the formula for the area of a triangle. \[A = \frac{ bh }{ 2 }\]

OpenStudy (ashontae19):

ok so is it D?

OpenStudy (ashontae19):

@Isaiah.Feynman

Vocaloid (vocaloid):

nope not D use the formula given area = (1/2)(x + 7)(3x) know where to go from here?

OpenStudy (isaiah.feynman):

I didn't solve it.

OpenStudy (ashontae19):

so it would it be 12 ?

Vocaloid (vocaloid):

please stop guessing

Vocaloid (vocaloid):

do you know how to multiply out area = (1/2)(x + 7)(3x)? it's ok to say no

OpenStudy (ashontae19):

yes

Vocaloid (vocaloid):

ok, let me know what you get

OpenStudy (ashontae19):

0.5x+3.5+2.5?

Vocaloid (vocaloid):

uh not quite

Vocaloid (vocaloid):

3x(x+7) = ?

Vocaloid (vocaloid):

hint: (use distribution)

OpenStudy (ashontae19):

3x+21

Vocaloid (vocaloid):

not quite what is 3x*x?

OpenStudy (ashontae19):

3x^2

Vocaloid (vocaloid):

good! now what is 3x*7?

OpenStudy (ashontae19):

21x

Vocaloid (vocaloid):

good! now put it all together what is 3x(x+7)?

OpenStudy (ashontae19):

3x^2+21x

Vocaloid (vocaloid):

good! now we have (1/2)(3x^2 + 21x) = 66 multiplying both sides by 2 gives us 3x^2 + 21x = 132 know where to go from here?

OpenStudy (ashontae19):

dont we divide?

Vocaloid (vocaloid):

what would you like to divide?

OpenStudy (ashontae19):

would we divide 132 and 66

Vocaloid (vocaloid):

nope, remember that 66 isn't part of our equation anymore our equation is now 3x^2 + 21x = 132

Vocaloid (vocaloid):

we CAN divide both sides by 3. what do we get then?

OpenStudy (ashontae19):

so we would divide 132 by 3?

Vocaloid (vocaloid):

yes, but also divide the left side too

Vocaloid (vocaloid):

what is 3x^2 + 21x divided by 3?

OpenStudy (ashontae19):

44

Vocaloid (vocaloid):

nope not quite

Vocaloid (vocaloid):

what is 3x^2 divided by 3?

OpenStudy (ashontae19):

3

Vocaloid (vocaloid):

nope...

Vocaloid (vocaloid):

hint: 3A divided by 3 is A 3x^2 divided by 3 is ...?

OpenStudy (ashontae19):

2

Vocaloid (vocaloid):

nope, follow the pattern please

OpenStudy (ashontae19):

x^2

Vocaloid (vocaloid):

good! now what is 21x divided by 3?

OpenStudy (ashontae19):

7x

Vocaloid (vocaloid):

good! now we divide the other side 132 divided by 3 = ?

OpenStudy (ashontae19):

44

Vocaloid (vocaloid):

good! now we have x^2 + 7x = 44 know what to do from here?

OpenStudy (ashontae19):

yea its 11

Vocaloid (vocaloid):

hold on a sec let me check my numbers

Vocaloid (vocaloid):

actually, no, x is not 11

Vocaloid (vocaloid):

do you know how to factor x^2 + 7x = 44?

Vocaloid (vocaloid):

unless you were saying that the height is 11?

OpenStudy (ashontae19):

yes thats supposed to be

Vocaloid (vocaloid):

ok, may i ask how you got 11 as your height? (i'm not saying you're wrong i just want to make sure you understand)

OpenStudy (ashontae19):

i finished your problem